# Foundation Models for Music: A Survey

### v1

@@ IV-H2 Long-Sequence Modelling
Approaches that modify the widely popular Transformer primarily focus on the two most important shortcomings of this architecture when handling long sequences: the complexity of attention and positional encoding. The quadratic complexity of the standard attention mechanism has been primarily tackled using sparsification in structured and unstructured forms leading to a drastic decrease in complexity with minimal or no loss in performance [CGRS19, DYY+19, GQL+19, YGG+19, BPC20, KKL20, QML+20, ZGD+20, XOG+22]. The standard (absolute) positional encoding seemingly appropriate for short sequences found in many speech and natural language processing tasks has been extended to various relative positional encoding schemes [SUV18, DYY+19, RSR+20b, WLQ+22] and hybrids of absolute and relative encodings [SAL+24, SDP+23]. Although many of these approaches have been integrated into Transformer-based foundational music models [HVU+18a, YLW+22], their ability to handle truly long music recordings remains to be seen. In this context the recent work on long-context speech recognition [FR23], where context length reached 1 hour of speech, suggests this to be a promising direction.
Recently there has also been substantial interest in developing novel approaches with linear or near-linear complexity with respect to sequence length. Among them structured state-space models [GGR22] appear to yield promising results for sequences reaching 1 million tokens [GD23]. These models can be interpreted as a combination of recurrent and convolutional architectures which additionally integrate principled mechanisms for modelling long-range dependencies. Another popular line of research explores hierarchical architectures where sequences are segmented into shorter units [DYY+19, PZV+19, RPJ+20]. This allows making use of separate mechanisms for modelling short-term (e.g. attention) dependencies within segments and long-term (e.g. implicit [BKKB24] or explicit [PZV+19] recurrence) dependencies across segments. Some of these approaches showed promising results on synthetic tasks with sequences reaching 2 million tokes.
@@ IV-H3 Causal Modelling
The discussion about domain knowledge earlier in this section stopped short of mentioning perhaps one of the strongest kinds of domain knowledge - causal relationships. Such relationships we cultivate all of our lives. In childhood, many of us learned the impact of key pair variations on our perception of sound by tinkering with the piano keys. As we grow older and start watching movies we begin to judge choices made by sound directors in particularly dramatic scenes. The knowledge of these relationships and our ability to enact them enables us to conduct interventions (e.g. fix a key, increase tempo) and create counterfactual pieces (e.g. imagine a known piece in a new key). Do machine learning engineers need to know the causal relationships that Horowitz formed and his ability to manipulate them in order to automatically generate Horowitz-quality piano pieces?
@@
Even in simpler scenarios than music determining causal relationships is a non-trivial endeavour. In many cases this is hardly a possible task due to the lack of access (e.g. astronomy) or understanding (e.g. neurology). In other cases, the full set of causal relationships is not known or available. Latent variables (e.g. hidden Markov models [Rab89], auto-encoders [Kra91]) emerged as a powerful but implicit surrogate aimed at capturing some of these relationships. However, the lack of appropriate parameter estimation methodologies capable of learning causal rather than merely correlational relationships meant that latent variable models have so far achieved limited success in this area. Causal learning emerged in response to these deficiencies of modern machine learning. Starting from the work done by Pearl and colleagues [Pea09] it has now emerged into a new field full of exciting developments [KLL+22]. Despite much progress made in causal learning over the years, work done in music and related areas remains extremely limited. The key issue stems from the lack of data with labelled causal information and/or methods capable of automatically extracting causal relationships. Although some causal models have emerged for simple image generation tasks, where causal factors could be limited to shape, texture and background [SG21], it is not clear how to extend them to other areas where causal relationships are much less trivial to specify, extract and model.
@@
An interesting alternative trend recently emerged where large non-causal models are manipulated into exhibiting causal relationships [LPP+20, MBAB22, OWJ+22, WWS+22, LPV+23]. One prominent example is the work of Wei and colleagues on chain-of-thought prompting for eliciting factually correct reasoning of large language models in question-answer type tasks [WWS+22]. Unlike standard prompting [BMR+20], where an example question is supplemented with a short factual answer (e.g. giving a numerical answer, such as 11, to the question about how many cans of food is left after a certain number of them goes missing and some restocked), the chain-of-thought prompting supplies an answer that showcases how a human could arrive to that conclusion [LYDB17]. This approach has been shown to be effective in arithmetic, common sense and reasoning tasks. Whether similar prompt manipulations can be extended to foundation music models remains to be seen.
@@ V-A Datasets
Data plays a central importance role in training the underlying model. The diversity, quantity and quality of the training data contribute greatly to the generalisability and robustness of the model[XPD+24, SKB+24, ZQL+24]. For machine learning and computer music researchers, training music-based models may require hundreds of billions of tokens or even larger datasets, and selecting high-quality datasets of different music recordings is a great challenge. In the realm of computer music, datasets are categorised into composition-level, typically sheets represented in MusicXML and ABC notation, performance-level symbolic music exemplified by MIDI formats, and datasets comprising raw audio waveforms. This section explores extensive open-source music datasets that are beneficial for training LLMs or LDMs tailored for music applications. Additionally, we will discuss Python libraries that are pivotal for music processing. Subsequently, the focus will shift to multimodal music datasets. Finally, the evaluation methodologies for foundation models developed on these datasets will be examined.
@@ V-A1 Composition-level Symbolic Music Datasets
In this subsection, we focus on introducing symbolic music datasets that contain more than 1,000 music excerpts. We will exclude smaller datasets though they are potentially useful for evaluation or supervised fine-tuning in specific domains. For more information of music datasets, readers can refer to the datasets website page of ISMIR community  [^15] https://www.ismir.net/resources/datasets/.
@@
Besides representing traditional Western Classical music score, MusicXML is widely utilised for storing lead sheets of popular music, encompassing the melody, lyrics, harmony, and various markings of a song. The melody is notated in modern Western music notation, with lyrics presented as text beneath the score, and harmony indicated by chord symbols positioned above the score. Notably, lead sheets provide little details regarding instrumentation or accompaniment.
The Wikifonia Lead Sheet Dataset [LRL17], originally from the now-defunct public repository Wikifonia.org, includes 5,533 MusicXML lead sheets of Western music across various genres. Before the service ceased in 2013, Lim et al. secured 2,252 lead sheets in major key with typically one chord per bar. If a bar contained multiple chords, only the first was chosen. The dataset is converted and distributed to CSV format on the website [^17] http://marg.snu.ac.kr/chord_generation/. Besides ABC notation, it also provide a version of MusicXML and MIDI files.
To complement the Yamaha Signature MIDI Collection, Jeong et al. developed the MuseScore Lead Sheet Dataset [JKK+19] by collecting MusicXML scores from MuseScore, a community-driven music score web platform, along with including their own transcriptions. These scores were transcribed voluntarily by community members, Although the dataset includes just 226 MusicXML pieces by 16 composers, it corresponds with 1,052 piano performances in MIDI format. Collectively, these pieces comprise a total of 666,918 notes in MusicXML and 3,547,683 notes in MIDI.
Hooktheory Lead Sheet Dataset (HLSD) [YHF+21] include 11,329 music lead sheet samples collected from the TheoryTab forum on Hooktheory’s website, a resource specializing in music education software. These samples whose melodies are transcribed by users with corresponding chord progressions in high-quality, denoted by both chord symbols and functional labels. The dataset included an excessive 704 chord class. Due to copyright restrictions, only song snippets are shared on TheoryTab, with annotations such as structural and genre labels.
@@
The ABC format is a plain text method for documenting music. Besides representing single-line melodies, it actually has the capability to fully represent polyphonic classical scores.
@@
Besides, there is other open-source website which has Western folk music in ABC format such as the Nottingham Music Dataset (NMD), comprises 1,200 British and American folk songs. A recently refined version is available online  [^20] https://ifdo.ca/seymour/nottingham/nottingham.html. Another example is the Henrik Norbeck collection that includes more than 2,800 ABC scores with lyrics, for Ireland and Sweden  [^21] http://www.norbeck.nu/abc/. More information can be found in the dataset section of the music generation survey [JLY20].
@@ V-A2 Performance-level Symbolic Music Datasets
The Lakh MIDI Dataset (LMD)[Raf16] is a large MIDI polyphonic music corpus with inconsistent expressive characteristics, consisting of various genres, instruments, and periods of time. The “LMD full” version includes the whole 176,581 MIDI files after removing duplication, and the “LMD matched” subset refers to the 45,129 files that aligned with items in the Million Song Dataset (MSD).
@@
The Augmented Design Lab (ADL) Piano MIDI dataset [FLW20] is derived from the LMD by retaining only one of the multiple versions of each song, and only remains tracks featuring piano-family instruments identified by MIDI program index 1-8, resulting in 9,021 unique rock or classical piano MIDI files. To enhance genre diversity with additional styles like jazz, they incorporated 2,065 files sourced from Internet resources, culminating in a total of 11,086 pieces in the final dataset.
@@
The NES-MDB dataset [DMM18] encompasses 5,278 songs from the soundtracks of 397 NES games, involving compositions from 296 composers and containing over two million notes. This dataset is extracted from the assembly code of NES games, capturing precise timings and parameters for authentic chiptune renditions. The NES synthesizer includes five instrument voices—two pulse-wave, one triangle-wave, and one noise generator—with the complexity of its audio synthesis chip abstracted for researchers. NES-MDB is available in multiple formats to facilitate research in NES music, with additional details provided for those interested in deeper technical exploration.
Besides, there are many online corpus for MIDI files such as BitMidi [^24] https://bitmidi.com/ Classical Archives [^25] https://www.classicalarchives.com/, and FreeMidi [^26] https://freemidi.org/.
@@
Apart from what has been mentioned above, there are multiple music transcription or generation datasets not only include symbolic music but audio as well.
MIDI and Audio Edited for Synchronous TRacks and Organization (MAESTRO) dataset [HSR+18] consists of about 200 hours of professional piano performances includeing both MIDI annotations and audio recordings.
@@ V-A3 Acoustic Music Datasets
MTG-Jamendo [BWT+19] Dataset is compiled from music available on Jamendo [^27] https://www.jamendo.com/, and features tags applied by the uploaders. This dataset encompasses 55,701 full audio tracks, each at least 30 seconds long and encoded at 320kbps MP3, totalling 509 GB of audio. The average track length is 224 seconds, summing up to 3,777 hours of audio. The tracks are annotated with 692 tags, derived from genres, instruments, and mood/theme categories. To streamline the dataset, variant spellings and semantically identical tags were consolidated, resulting in 99 remapped tags.
@@
Besides, there are many large-scale datasets which only provide the feature or URL of the audio, allowing users to collect the tracks from the internet, for example Million Song Dataset (MSD) [BMEWL11], AudioSet [GEF+17], AcousticBrainz [PBK+15], and Disco-10M [LGFW24] etc.
@@ V-A5 Multimodal Datasets: Present & Challenges
MusicQA [LHSS24] is a dataset of music question-answer pairs, generated by MPT-7B model [Tea23]. The datase is create from MusicCaps and MagnaTagATune datasets that are music captions or tags related to descriptive and inferential questions about the music. Each audio has five open-ended question-answer pairs focusing on aspects like music emotion, tempo, and genre. It include 12,542 music clips in the training set, 76.15 hours in total, with 112,878 question-answer pairs.
MusicInstruct [DML+24], also derived from the MusicCaps dataset’s music-caption pairs, contains Q&A pairs generated with few-shot learning techniques by GPT-4. It includes 5,521 YouTube ID of audio clips; a short version with 27,540 Q&A pairs focused on detailed aspects such as emotion, instruments, vocals, tempo, and genre, typically eliciting concise one or two-sentence responses; as well as a long Version that contains 32,953 Q&A pairs with broader questions about musical pieces, often resulting in more elaborate, paraphrased responses.
@@
The MUEdit dataset [HLSS23a] is curated to facilitate the development and training of models capable of understanding and editing music based on specific prompts. The MUEdit dataset includes 10-second music pairs, each selected based on rhythmic characteristics such as tempo and pitch to ensure consistency and relevancy, 55.69 hours in total. The creation process involves generating descriptions for these music pairs using the MU-LLaMA model, followed by the generation of editing instructions via the MPT-7B model. These instructions are designed to mimic a conversational interface where the model not only receives detailed descriptions but also generates editing commands based on these descriptions, simulating a realistic user-model interaction for music editing tasks.
WikiMusicText (WikiMT) [WYTS23a] contains 1010 lead sheets in ABC notation format from Wikifonia.org, each with metadata including title and artist extracted directly from the scores. Besides, it also derives genre and description from Internet information. The genre labels are assigned by analysing related Wikipedia content and categorising it into 8 GTZAN taxonomy classes: Jazz, Country, Folk, R&B, Pop, Rock, Dance, and Latin. Descriptions are synthesised from Wikipedia using the BART-large model. To emphasise musical content, extraneous text within the ABC notation has been removed.
@@
Human experts conduct the annotation process for EmoMV-A, while a pre-trained deep neural network generates annotations for EmoMV-B and EmoMV-C. The reliability of these emotional labels is evaluated through a user study that examines their accuracy. In conjunction with the dataset development, researchers have created a benchmark deep learning model. This model is designed for binary classification of affective correspondence between music and video, and has been optimised for affective music-video retrieval applications. This advancement provides a valuable resource for investigating how music and visual elements interact to express emotional content.
SymMV dataset [ZWW+23b] is a larger video-music dataset that pairs music videos (MV) with piano MIDI and audio. It includes 1140 pop piano music pieces, along with their corresponding official MV, totalling 76.5 hours. Each MV pair is annotated for chord progression, tonality, and rhythm, enhancing its utility for research in MV emotional correspondence. The dataset functions as a valuable resource for developing algorithms that generate background music for videos. It enables the exploration of inherent relationships between visual and musical elements, leveraging the emotional and stylistic correspondence between video content and its soundtrack. The creation of this dataset involved a meticulous process: sourcing professional-grade piano renditions from select YouTube channels, aligning these with the original music videos, and utilizing sophisticated transcription algorithms to convert the audio into MIDI format. To maintain high standards, professional musicians conducted a thorough review of the compiled dataset.

### v2

@@ IV-H2 Long-Sequence Modelling
Approaches that modify the widely popular Transformer primarily focus on the two most important shortcomings of this architecture when handling long sequences: the complexity of attention and positional encoding. The quadratic complexity of the standard attention mechanism has been primarily tackled using sparsification in structured and unstructured forms leading to a drastic decrease in complexity with minimal or no loss in performance [CGRS19, DYY+19, GQL+19, YGG+19, BPC20, KKL20, QML+20, ZGD+20, XOG+22]. The standard (absolute) positional encoding seemingly appropriate for short sequences found in many speech and natural language processing tasks has been extended to various relative positional encoding schemes [SUV18, DYY+19, RSR+20b, WLQ+22] and hybrids of absolute and relative encodings [SAL+24, SDP+23]. Although many of these approaches have been integrated into Transformer-based foundational music models [HVU+18a, YLW+22], their ability to handle truly long music recordings remains to be seen. In this context, the recent work on long-context speech recognition [FR23], where context length reached 1 hour of speech, suggests this to be a promising direction.
Recently there has also been substantial interest in developing novel approaches with linear or near-linear complexity with respect to sequence length. Among them, structured state-space models [GGR22] appear to yield promising results for sequences reaching 1 million tokens [GD23]. These models can be interpreted as a combination of recurrent and convolutional architectures which additionally integrate principled mechanisms for modelling long-range dependencies. Another popular line of research explores hierarchical architectures where sequences are segmented into shorter units [DYY+19, PZV+19, RPJ+20]. This allows making use of separate mechanisms for modelling short-term (e.g. attention) dependencies within segments and long-term (e.g. implicit [BKKB24] or explicit [PZV+19] recurrence) dependencies across segments. Some of these approaches showed promising results on synthetic tasks with sequences reaching 2 million tokes.
@@ IV-H3 Causal Modelling
The discussion about domain knowledge earlier in this section stopped short of mentioning perhaps one of the strongest kinds of domain knowledge - causal relationships. Such relationships we cultivate all of our lives. In childhood, many of us learned the impact of key pair variations on our perception of sound by tinkering with the piano keys. As we grow older and start watching movies we begin to judge choices made by sound directors in particularly dramatic scenes. The knowledge of these relationships and our ability to enact them enables us to conduct interventions (e.g. fix a key, increase tempo) and create counterfactual pieces (e.g. imagine a known piece in a new key). Do machine learning engineers need to know the causal relationships that the pianist Horowitz formed and his ability to manipulate them in order to automatically generate Horowitz-quality piano pieces?
@@
Even in simpler scenarios than music, determining causal relationships is a non-trivial endeavour. In many cases this is hardly a possible task due to the lack of access (e.g. astronomy) or understanding (e.g. neurology). In other cases, the full set of causal relationships is not known or available. Latent variables (e.g. hidden Markov models [Rab89], auto-encoders [Kra91]) emerged as a powerful but implicit surrogate aimed at capturing some of these relationships. However, the lack of appropriate parameter estimation methodologies capable of learning causal rather than merely correlational relationships meant that latent variable models have so far achieved limited success in this area. Causal learning emerged in response to these deficiencies of modern machine learning. Starting from the work done by Pearl and colleagues [Pea09] it has now emerged into a new field full of exciting developments [KLL+22]. Despite much progress made in causal learning over the years, work done in music and related areas remains extremely limited. The key issue stems from the lack of data with labelled causal information and/or methods capable of automatically extracting causal relationships. Although some causal models have emerged for simple image generation tasks, where causal factors could be limited to shape, texture and background [SG21], it is not clear how to extend them to other areas where causal relationships are much less trivial to specify, extract, and model.
@@
An interesting alternative trend recently emerged where large non-causal models are manipulated into exhibiting causal relationships [LPP+20, MBAB22, OWJ+22, WWS+22, LPV+23]. One prominent example is the work of Wei and colleagues on chain-of-thought prompting for eliciting factually correct reasoning of large language models in question-answer type tasks [WWS+22]. Unlike standard prompting [BMR+20], where an example question is supplemented with a short factual answer (e.g. giving a numerical answer, such as 11, to the question about how many cans of food are left after a certain number of them go missing and some restocked), the chain-of-thought prompting supplies an answer that showcases how a human could arrive to that conclusion [LYDB17]. This approach has been shown to be effective in arithmetic, common sense and reasoning tasks. Whether similar prompt manipulations can be extended to music foundation models remains to be seen.
@@ V-A Datasets
Data plays a central role in training the underlying model. The diversity, quantity and quality of the training data contribute greatly to the generalisability and robustness of the model[XPD+24, SKB+24, ZQL+24]. For machine learning and computer music researchers, training music-based models may require hundreds of billions of tokens or even larger datasets, and selecting high-quality datasets of different music recordings is a great challenge. In the realm of computer music, datasets are categorised into composition-level, typically sheets represented in MusicXML and ABC notation, performance-level symbolic music exemplified by MIDI formats, and datasets comprising raw audio waveforms. This section explores extensive open-source music datasets that are beneficial for training LLMs or LDMs tailored for music applications. Additionally, we will discuss Python libraries that are pivotal for music processing. Subsequently, the focus will shift to multimodal music datasets. Finally, the evaluation methodologies for foundation models developed on these datasets will be examined.
@@ V-A1 Composition-level Symbolic Music Datasets
In this subsection, we focus on introducing symbolic music datasets that contain more than 1,000 music excerpts. We will exclude smaller datasets though they are potentially useful for evaluation or supervised fine-tuning in specific domains. For more information on music datasets, readers can refer to the datasets website page of the ISMIR community  [^15] https://www.ismir.net/resources/datasets/.
@@
Besides representing traditional Western Classical music scores, MusicXML is widely utilised for storing lead sheets of popular music, encompassing the melody, lyrics, harmony, and various markings of a song. The melody is notated in modern Western music notation, with lyrics presented as text beneath the score, and harmony indicated by chord symbols positioned above the score. Notably, lead sheets provide little details regarding instrumentation or accompaniment.
The Wikifonia Lead Sheet Dataset [LRL17], originally from the now-defunct public repository Wikifonia.org, includes 5,533 MusicXML lead sheets of Western music across various genres. Before the service ceased in 2013, Lim et al. secured 2,252 lead sheets in major keys with typically one chord per bar. If a bar contained multiple chords, only the first was chosen. The dataset is converted and distributed to CSV format on the website [^17] http://marg.snu.ac.kr/chord_generation/. Besides ABC notation, it also provides a version of MusicXML and MIDI files.
To complement the Yamaha Signature MIDI Collection, Jeong et al. developed the MuseScore Lead Sheet Dataset [JKK+19] by collecting MusicXML scores from MuseScore, a community-driven music score web platform, along with including their own transcriptions. These scores were transcribed voluntarily by community members. Although the dataset includes just 226 MusicXML pieces by 16 composers, it corresponds to 1,052 piano performances in MIDI format. Collectively, these pieces comprise a total of 666,918 notes in MusicXML and 3,547,683 notes in MIDI.
Hooktheory Lead Sheet Dataset (HLSD) [YHF+21] include 11,329 music lead sheet samples collected from the TheoryTab forum on Hooktheory’s website, a resource specializing in music education software. These samples, whose melodies are transcribed by users with corresponding chord progressions in high quality, are denoted by both chord symbols and functional labels. The dataset included an excessive 704 chord class. Due to copyright restrictions, only song snippets are shared on TheoryTab, with annotations such as structural and genre labels.
@@
The ABC format is a plain text method for documenting music. Besides representing single-line melodies, it has the capability to fully represent polyphonic classical scores.
@@
Further, an open-source website has Western folk music in ABC format such as the Nottingham Music Dataset (NMD), which is comprised of 1,200 British and American folk songs. A recently refined version is available online  [^20] https://ifdo.ca/seymour/nottingham/nottingham.html. Another example is the Henrik Norbeck collection that includes more than 2,800 ABC scores with lyrics, for Ireland and Sweden  [^21] http://www.norbeck.nu/abc/. More information can be found in the dataset section of the music generation survey [JLY20].
@@ V-A2 Performance-level Symbolic Music Datasets
The Lakh MIDI Dataset (LMD)[Raf16] is a large MIDI polyphonic music corpus with inconsistent expressive characteristics, consisting of various genres, instruments, and periods of time. The “LMD full” version includes the whole 176,581 MIDI files after removing duplication, and the “LMD matched” subset refers to the 45,129 files that are aligned with items in the Million Song Dataset (MSD).
@@
The Augmented Design Lab (ADL) Piano MIDI dataset [FLW20] is derived from the LMD by retaining only one of the multiple versions of each song and only tracks featuring piano-family instruments identified by MIDI program index 1-8, resulting in 9,021 unique rock or classical piano MIDI files. To enhance genre diversity with additional styles like jazz, they incorporated 2,065 files sourced from Internet sources, culminating in a total of 11,086 pieces in the final dataset.
@@
The NES-MDB dataset [DMM18] encompasses 5,278 songs from the soundtracks of 397 NES games, involving compositions from 296 composers and containing over two million notes. This dataset is extracted from the assembly code of NES games, capturing precise timings and parameters for authentic chiptune renditions. The NES synthesiser includes five instrument voices—two pulse-wave, one triangle-wave, and one noise generator—with the complexity of its audio synthesis chip abstracted for researchers. NES-MDB is available in multiple formats to facilitate research in NES music, with additional details provided for those interested in deeper technical exploration.
Additionally, there are many online corpora for MIDI files such as BitMidi [^24] https://bitmidi.com/ Classical Archives [^25] https://www.classicalarchives.com/, and FreeMidi [^26] https://freemidi.org/.
@@
Apart from what has been mentioned above, there are multiple music transcription or generation datasets that not only include symbolic music but audio as well.
MIDI and Audio Edited for Synchronous TRacks and Organization (MAESTRO) dataset [HSR+18] consists of about 200 hours of professional piano performances including both MIDI annotations and audio recordings.
@@ V-A3 Acoustic Music Datasets
The MTG-Jamendo [BWT+19] Dataset is compiled from music available on Jamendo [^27] https://www.jamendo.com/, and features tags applied by the uploaders. This dataset contains 55,701 full audio tracks, each at least 30 seconds long and encoded at 320kbps MP3, totalling 509 GB of audio. The average track length is 224 seconds, summing up to 3,777 hours of audio. The tracks are annotated with 692 tags, derived from genres, instruments, and mood/theme categories. To streamline the dataset, variant spellings and semantically identical tags were consolidated, resulting in 99 remapped tags.
@@
Aside from the aforementioned, there are many large-scale datasets which only provide the feature or URL of the audio, allowing users to collect the tracks from the internet. Some such examples include the Million Song Dataset (MSD) [BMEWL11], AudioSet [GEF+17], AcousticBrainz [PBK+15], and Disco-10M [LGFW24] etc.
@@ V-A5 Multimodal Datasets: Present & Challenges
MusicQA [LHSS24] is a dataset of music question-answer pairs, generated by MPT-7B model [Tea23]. The dataset is created by using MusicCaps and MagnaTagATune that contain music captions or tags related to descriptive and inferential questions about the music. Each audio has five open-ended question-answer pairs focusing on aspects like music emotion, tempo, and genre. It include 12,542 music clips in the training set, 76.15 hours in total, with 112,878 question-answer pairs.
MusicInstruct [DML+24], also derived from the MusicCaps dataset’s music-caption pairs, contains Q&A pairs generated with few-shot learning techniques by GPT-4. It includes 5,521 YouTube ID of audio clips, a short version with 27,540 Q&A pairs focused on detailed aspects such as emotion, instruments, vocals, tempo, and genre, typically eliciting concise one or two-sentence responses, as well as a long Version that contains 32,953 Q&A pairs with broader questions about musical pieces, often resulting in more elaborate, paraphrased responses.
@@
The MUEdit dataset [HLSS23a] is curated to facilitate the development and training of models capable of understanding and editing music based on specific prompts. The MUEdit dataset includes 10-second music pairs, each selected based on rhythmic characteristics such as tempo and pitch to ensure consistency and relevancy, totalling 55.69 hours. The creation process involves generating descriptions for these music pairs using the MU-LLaMA model, followed by the generation of editing instructions via the MPT-7B model. These instructions are designed to mimic a conversational interface where the model not only receives detailed descriptions but also generates editing commands based on these descriptions, simulating a realistic user-model interaction for music editing tasks.
WikiMusicText (WikiMT) [WYTS23a] contains 1010 lead sheets in ABC notation format from Wikifonia.org, each with metadata including title and artist extracted directly from the scores. Additionally, it derives genre and description from Internet information. The genre labels are assigned by analysing related Wikipedia content and categorising it into 8 GTZAN taxonomy classes: Jazz, Country, Folk, R&B, Pop, Rock, Dance, and Latin. Descriptions are synthesised from Wikipedia using the BART-large model. To emphasise musical content, extraneous text within the ABC notation has been removed.
@@
Human experts conduct the annotation process for EmoMV-A, while a pre-trained deep neural network generates annotations for EmoMV-B and EmoMV-C. The reliability of these emotional labels is evaluated through a user study that examines their accuracy. In conjunction with the dataset development, researchers have created a benchmark deep learning model. This model is designed for binary classification of affective correspondence between music and video and has been optimised for affective music-video retrieval applications. This advancement provides a valuable resource for investigating how music and visual elements interact to express emotional content.
SymMV dataset [ZWW+23b] is a larger video-music dataset that pairs music videos (MV) with piano MIDI and audio. It includes 1140 pop piano music pieces, along with their corresponding official MV, totalling 76.5 hours. Each MV pair is annotated for chord progression, tonality, and rhythm, enhancing its utility for research in MV emotional correspondence. The dataset functions as a valuable resource for developing algorithms that generate background music for videos. It enables the exploration of inherent relationships between visual and musical elements, leveraging the emotional and stylistic correspondence between video content and its soundtrack. The creation of this dataset involved a meticulous process: sourcing professional-grade piano renditions from select YouTube channels, aligning these with the original music videos, and utilising sophisticated transcription algorithms to convert the audio into MIDI format. To maintain high standards, professional musicians conducted a thorough review of the compiled dataset.